Weights of 5, 6, 8, and 10 N are hung from a 12 m rod at distances 2, 5, 7, and 8 m from the left end. Neglecting theweight of the rod, how far from the left end is the center of gravity?
1
Expert's answer
2019-06-21T08:25:11-0400

For coordinates center of gravity

"x_c=\frac{x_1 m_1+x_2 m_2+x_3 m_3+x_4 m_4}{m_1+m_2+m_3+m_4}"

"x_c=\frac{2\cdot 5+5\cdot 6+7\cdot 8+8\cdot 10}{5+6+8+10}=6.1 (m)"